# A Simple Lifelong Learning Approach

### v19

@@ 3 Representation ensembling algorithms
A canonical example of a single learner depicted in Figure 2A is a decision tree. Importantly, one can subsample the training data to learn different components of the tree [33, 34, 35]. For example, one can use a portion of data to learn the tree structure (which is the encoder). Then, by pushing the remaining data (sometimes called the ‘out-of-bag’ data) through the tree, one can learn posteriors in each leaf node (which are the channel). The channel thus gives scores for each data point denoting the probability of that data point belonging to a specific class. Using separate sets of data to learn the encoder and the channel results in less bias in the estimated posterior in the channels as in ‘honest trees’ [33, 34, 35]. Finally, the decoder provides the predicted class label using $\operatornamewithlimits{argmax}$ over the posteriors from the channel.
@@
3.1 Our representation ensembling algorithms
@@
3.1.1 Simple Lifelong Learning Networks in resource growing mode
A Simple Lifelong Learning Network (SiLLy-N) ensembles deep networks. For each task, the encoder $u_{t}$ in SiLLy-N is the “backbone” of a deep network (DN). Thus, each $u_{t}$ maps an element of $\mathcal{X}$ to an element of $\mathbb{R}^{d}$, where $d$ is the number of neurons in the penultimate layer of the DN. The channels are learned by averaging over $k$-Nearest Neighbors ($k$-NN) [39] trained over the $d$ dimensional representations of $\mathcal{X}$. Note that the channel is trained on the $d$ dimensional outputs from the encoders which is much smaller in size than the original training data and hence, the $k$-NN channels are inexpensive storage-wise (shown later in Figure 3). Other algorithms could also be used to learn the channels, though we do not pursue them here. The decoder $w_{t}$ outputs the $\operatornamewithlimits{argmax}$ to produce a single prediction.
3.1.2 Simple Lifelong Learning Networks in resource constrained mode
@@
3.1.3 Additional realization of our approach using random forest as encoder
Simple Lifelong Learning Forest (SiLLy-F) ensembles decision trees or forests. For each task, the encoder $u_{t}$ of SiLLy-F is the representation learned by a decision forest [40, 38]. The channel then learns the class-conditional posteriors by populating the forest leaves with out-of-task samples, as in “honest trees” [33, 34, 35]. Each channel outputs the posteriors averaged across the collection of forests learned over different tasks. The decoder $w_{t}$ outputs the argmax to produce a single prediction.
@@ 3.1.3 Additional realization of our approach using random forest as encoder
4 A computational taxonomy of lifelong learners
Figure 3: Storage space as a function on number of tasks in CIFAR 10X10. Memory consumed by SiLLy-N is dominated by the encoder size. The size of DF-CNN remains constant throughout.
The space complexity of the learner refers to the amount of memory space needed to store the learner [41]. We also study the representation capacity of these algorithms. Capacity is defined as the size of the subset of hypotheses that is achievable by the learning algorithm [42].
  1. 1.
    Each task has the same number of training samples.
  2. 2.
  3. 3.
    The number of epochs is fixed for each task.
  4. 4.
Table 1: Capacity, space, and time complexity of the representation learned by various lifelong learning algorithms. We show soft-O notation ($\tilde{\mathcal{O}}(\cdot,\cdot)$ defined in main text) as a function of $n=\sum_{t}^{T}n_{t}$ and $T$, as well as the common setting where $n$ is proportional to $T$. The bottom three rows show algorithms whose space and time both grow quasilinearly with capacity growing.
Parametric Capacity Space Time Examples
($n,T$) ($n,T$) ($n\propto T$) $(n,T)$ $(n\propto T)$
parametric $1$ $1$ $1$ $1$ $1$ SiLLy-N-M, SiLLy-F-M
parametric $1$ $1$ $1$ $n$ $n$ O-EWC, SI, LwF
parametric $1$ $T$ $n$ $nT$ $n^{2}$ EWC
parametric $1$ $n$ $n$ $nT$ $n^{2}$ Total Replay
semi-parametric $T$ $T^{2}$ $n^{2}$ $nT$ $n^{2}$ ProgNN
semi-parametric $T$ $T$ $n$ $n$ $n$ DF-CNN
semi-parametric $T$ $T+n$ $n$ $n$ $n$ SiLLy-N, Model Zoo, DER, LMC
non-parametric $n$ $n$ $n$ $n$ $n$ SiLLy-F, IBP-WF
Assumption $3$ enables us to write time complexity as a function of the sample size. Table 1 summarizes the capacity, space and time complexity of several reference algorithms, as well as our SiLLy-N and SiLLy-F. For space and time complexity, the table shows results as a function of $n$ and $T$, as well as the common scenario where sample size is fixed per task, and therefore proportional to the number of tasks, $n\propto T$. For detailed calculation of time complexity see Appendix A.3.
Parametric lifelong learning methods have a representational capacity which is invariant to sample size and task number. Although the space complexity of some of these algorithms grow (because the number of the constraints stored by the algorithms grows, or they continue to store more data), their capacity is fixed. Thus, given a sufficiently large number of tasks with increasing complexity, in general, eventually all parametric methods will catastrophically forget. EWC [10], Online EWC [13], SI [11], and LwF [12] are all examples of parametric lifelong learning algorithms. Our fixed resource algorithms are also parametric. For comparison, we implement another baseline algorithm and refer to it as Total Replay, which is also parametric. Total Replay replays both old and current task data while learning a new task.
Semi-parametric algorithms’ representational capacity grows slower than sample size. For example, if $T$ is increasing slower than $n$ (e.g., $T\propto\log n$), then algorithms whose capacity is proportional to $T$ are semi-parametric. ProgNN [16] is semi-parametric, nonetheless, its space complexity is $\tilde{\mathcal{O}}(T^{2})$ due to the lateral connections. Moreover, the time complexity for ProgNN also scales quadratically with $n$ when $n\propto T$. Thus, an algorithm that literally stores all the data it has ever seen, and retrains a fixed size network on all those data with the arrival of each new task, would have smaller space complexity and the same time complexity as ProgNN. DF-CNN [17] improves upon ProgNN by introducing a “knowledge base” with lateral connections to each new column, thereby avoiding all pairwise connections. Because these semi-parametric methods have a fixed representational capacity per task, they will either lack the representation capacity to perform well given sufficiently complex tasks, and/or will waste resources for very simple tasks.
SiLLy-N and SiLLy-F eliminate the lateral connections between columns of the network, thereby reducing space complexity down to $\tilde{\mathcal{O}}(T)$. Moreover, as shown in Figure 3, memory consumed by new channels is negligible compared to that of memory required for storing the encoders. Note that the time required for updating channels is negligible in comparison with the time required for training a new encoder.
Indian Buffet Process for Weight Factors (IBP-WF) is another non-parametric lifelong learning algorithm.
5 Providing intuition of simple lifelong learning through simulations
@@
5.1 Forward and backward transfer in a simple environment
Consider a very simple two-task environment: Gaussian XOR and Gaussian Exclusive NOR (XNOR) (Figure 4A, see Appendix A.4 for details). The two tasks share the exact same discriminant boundaries: the coordinate axes. Thus, transferring from one task to the other merely requires learning a bit flip of the class labels. We sample a total 750 samples from XOR, followed by another 750 samples from XNOR.
@@ 5.1 Forward and backward transfer in a simple environment
5.2 Forward and backward transfer for adversarial tasks
@@
6 Benchmark data experiments
For benchmark data, we build SiLLy-N encoders using the network architecture described in [45]. We use the same network architecture for all the benchmarking models. For the following experiments, we consider two modalities of real data: vision and language.
6.1 Reference algorithms
We compared our approaches to $15$ reference lifelong learning methods. Among them five are resource growing as well as modular approach: ProgNN [16], DF-CNN [17], LMC [46], Model Zoo[47], CoSCL [48]. Note that "Model Zoo" was published after our work was archived on arXiv, and the authors have built on our work (personal communications). Other reference algorithms are resource constrained: Elastic Weight Consolidation (EWC) [10], Online-EWC (O-EWC) [13], Synaptic Intelligence (SI) [11], Learning without Forgetting (LwF) [12], and “None”.
We also compare two variants of exact replay (Total Replay and Partial Replay) using the code provided in [45]. Both Total and Partial Replay store all the data they have ever seen, but Total Replay replays all of it upon acquiring a new task, whereas Partial Replay replays $N$ samples, randomly sampled from the entire corpus, whenever we acquire a new task with $N$ samples. Additionally, we have compared our approach with more constrained ways of replaying old task data, including Averaged Gradient Episodic Memory (A-GEM) [49], Experience Replay (ER) [50] and Task-based Accumulated Gradients (TAG) [51].
For the baseline “None”, the network was incrementally trained on all tasks in the standard way while always only using the data from the current task. The implementations for all of the algorithms are adapted from open source codes [17, 52]; for implementation details, see Appendix A.2.
6.2 Exploring and explaining transfer capabilities via CIFAR 10x10 dataset
The CIFAR 100 challenge [53], consists of 50,000 training and 10,000 test samples, each a 32x32 RGB image of a common object, from one of 100 possible classes, such as apples and bicycles. CIFAR 10x10 divides these data into 10 tasks, each with 10 classes [17] (see Appendix A.5 for details).
6.2.1 Resource growing experiments
SiLLy-N and Model Zoo demonstrate positive forward and backward transfer for every task in CIFAR 10x10, in contrast, other algorithms do not exhibit any positive backward transfer (Figure 1 first column). Moreover, they retained their accuracy while improving transfer (Figure 1, bottom row). ProgNN had a similar degree of forward transfer, but zero backward transfer, and requires quadratic space and time in sample size, unlike SiLLy-N which requires quasilinear space and time.
@@ 6.2.1 Resource growing experiments
6.2.2 Ablation experiments
Our proposed algorithms can improve performance on all the tasks (past and future) by both growing additional resources and replaying data from the past tasks. Below we do two ablation experiments using CIFAR 10X10 to measure the relative contribution of resource growth and replay on the performance of our proposed algorithms.
@@ 6.2.2 Ablation experiments
6.2.3 Experiment using pretrained encoders
Figure 6: Pretrained encoder on CIFAR 10X10. Using pretrained encoders results in better transfer and accuracy for SiLLy-N.
In this experiment, we explore the effect of using pretrained encoders on the performance of SiLLy-N. For this experiment only, we use ResNet 50, as weights for the model pretrained on imagenet dataset is publicly available with the python package: Keras [54]. We initialize the weights of the encoders with the pretrained weights and do not freeze any layer during training. As shown in Figure 6, pretraining the encoders results in better accuracy and forward transfer, but less backward transfer for SiLLy-N. However, as it is unclear how to use pretrained encoders for other baseline approaches, we do not use pretrained encoders for SiLLy-N in other experiments for fair comparison.
6.2.4 Adversarial experiments
Consider the same CIFAR 10x10 experiments above, but, for Task 2 through 9, randomly permute the class labels within each task, rendering each of those tasks adversarial with regard to the first task (because the labels are uninformative). Figure 7A indicates that backward transfer for SiLLy-N show positive backward transfer even with such label shuffling (the other algorithms did not demonstrate positive backward transfer). Now, consider a Rotated CIFAR experiment, which uses only data from the first task, divided into two equally sized subsets (making two tasks), where the second subset is rotated by different amounts (Figure 7, right). Backward transfer of SiLLy-N is nearly invariant to rotation angle, whereas the other approaches are far more sensitive to rotation angle. Note that zero rotation angle corresponds to the two tasks having identical distributions. The fact that other algorithms fail to transfer even in this setting suggests that they may not ever be able to positively backwards transfer. See Appendix A.5 for additional experiment using CIFAR 10X10.
@@ 6.2.4 Adversarial experiments
Figure 7: Extended CIFAR 10x10 experiments. A. Shuffling class labels within tasks two through nine with 500 samples each demonstrates SiLLy-N can still achieve positive backward transfer, and that the other algorithms still fail to transfer. B. SiLLy-N is nearly invariant to rotations, whereas other approaches are more sensitive to rotation.
6.3 Further investigating transfer in additional datasets with more classes, tasks, and/or samples
Figure 8: Performance summary on vision and speech benchmark datasets with varying sample sizes per task. Dataset on the right has more samples per task. Imagenet and five dataset have really high sample size per task which is a non-trivial lifelong learning setting. SiLLy-N performs the best in the low sample size regime on the left two columns which is desirable in lifelong learning. See appendix Figure 3 and 5 for the extended results.
6.3.1 Spoken Digit
In this experiment, we used the Spoken Digit dataset [55]. As shown in Figure 8 first column, SiLLy-N shows positive backward and forward transfer between the spoken digit tasks, in contrast to other methods, some of which show only forward transfer, others show only backward transfer, with none showing both, and some showing neither. See Appendix A.5 for details of the experiment.
6.3.2 FOOD1k 50X20 Dataset
In this experiment, we use Food1k which is a large scale vision dataset consisting of $1000$ food categories from Food2k [56]. FOOD1k 50X20 splits these data into $50$ tasks with $20$ classes each. For each class, we randomly sampled $60$ samples per class for training the models and used rest of the data for testing purpose. Because on the CIFAR experiments Model Zooperforms the best among the reference resource growing models, and LwF is the best performing resource constrained algorithm, we only use them as the reference models for the large scale experiment to avoid heavy computational cost. As shown in Figure 8 second column, SiLLy-N performs the best among all the algorithms on this large dataset.
In lifelong learning, we are often primarily concerned with situations in which we have a small number of samples per task. If we have enough samples per tasks, the learning agent need not transfer knowledge from other tasks. However, below we also experiment with non-trivial lifelong learning setting where sample per task is high.
6.3.3 Split Mini-Imagenet
In this experiment, we have used the Mini-Imagenet dataset [51]. The dataset was split into $20$ tasks with $5$ classes each. Each task has $2400$ training samples and $600$ testing samples. As shown in Figure 1 fourth column, we get positive forward and backward transfer for SiLLy-N. However, although samples per task is lower compared to that of 5-dataset, it is still quite high. Hence, Model Zoo outperforms all the algorithms in this experiment.
6.3.4 5-dataset
In this experiment, we have used 5-dataset [51]. It consists of $5$ tasks from five different datasets: CIFAR-10 [53], MNIST, SVHN [57], notMNIST [58], Fashion-MNIST [59]. All the monochromatic images were converted to RGB format, and then resized to $3\times 32\times 32$. As shown in Appendix Table 3, training samples per task in $5$-dataset is relatively higher than that of low data regime typically considered in lifelong learning setting. However, as shown in Figure 8 fourth column, SiLLy-N show less forgetting than most of the reference algorithms. On the other hand, Model Zoo shows comparatively better performance in relatively high task data size setup. Recall that SiLLy-N is based on bagging, and Model Zoo is based on boosting. It is well known that boosting often outperforms bagging when sample sizes are large  [^2] Authors in [60] shows that both bagging and boosting asymptotically converge to the Bayes optimal solution. However, for finite sample size and similar model complexity, we empirically find bagging approach to lifelong learning performs better than that of boosting when the training sample size is low whereas boosting performs better on large training sample size (See Figure 1 and 8). This is consistent with similar results in single task learning [61, 62, 25].
6.4 Constant Resource Mode Operation

### v21

@@ IV Representation ensembling algorithms
A canonical example of a single learner depicted in Figure 2A is a decision tree. Importantly, one can subsample the training data to learn different components of the tree [53, 54, 55]. For example, one can use a portion of data to learn the tree structure (which is the encoder). Then, by pushing the remaining data (sometimes called the ‘out-of-bag’ data) through the tree, one can learn posteriors in each leaf node (which are the channel). The channel thus gives scores for each data point denoting the probability of that data point belonging to a specific class. Using separate sets of data to learn the encoder and the channel results in less bias in the estimated posterior in the channels as in ‘honest trees’ [53, 54, 55]. Finally, the decoder provides the predicted class label using $\operatorname*{arg\,max}$ over the posteriors from the channel.
@@
IV-A Our representation ensembling algorithms
Figure 3: Space and time complexity as a function on number of tasks in CIFAR 10X10. Section VI describes the experimental setup and architecture of the convolutional network used. Top left: Model size, SiLLy can flexibly operate between resource constrained and growing modes. Top right: Memory consumed by SiLLy-N is dominated by the encoder size. Bottom left: SiLLy takes less training time compared to other baselines. Bottom right: Inference time taken by different lifelong learners for $1000$ samples. Bottom row has log y-axis.
@@ IV-A Our representation ensembling algorithms
IV-A1 Simple Lifelong Learning Networks in resource growing mode
A Simple Lifelong Learning Network (SiLLy-N) ensembles deep networks. For each task, the encoder $u_{t}$ in SiLLy-N is the “backbone” of a deep network (DN). Thus, each $u_{t}$ maps an element of $\mathcal{X}$ to an element of $\mathbb{R}^{d}$, where $d$ is the number of neurons in the penultimate layer of DN. The channels are learned by averaging the outputs from decision forests [59, 58] trained on the $d$ dimensional representations of $\mathcal{X}$. See Appendix Figure 3 where we conduct experiments on a CIFAR 10X10 (described later in Section VI-B1) while varying the number of decision trees per channel. Appendix Figure 3 shows that decision forest-based channels are robust to hyperparameter perturbation. Other algorithms could also be used to learn the channels, though we do not pursue them here. The decoder $w_{t}$ outputs the $\operatorname*{arg\,max}$ to produce a single prediction.
IV-A2 Simple Lifelong Learning Networks in resource constrained mode
@@
IV-A3 Additional realization of our approach using random forest as encoder
Simple Lifelong Learning Forest (SiLLy-F) ensembles decision trees or forests. For each task, the encoder $u_{t}$ of SiLLy-F is the representation learned by a decision forest. The channel then learns the class-conditional posteriors by populating the forest leaves with out-of-task samples, as in “honest trees” [53, 54, 55]. Each channel outputs the posteriors averaged across the collection of forests learned over different tasks. The decoder $w_{t}$ outputs the argmax to produce a single prediction.
@@ IV-A3 Additional realization of our approach using random forest as encoder
Figure 3 top left shows the model size for our proposed approach grows linearly with the number of tasks. Moreover, the memory consumed by the new channels is negligible compared to the memory required to store the encoders (Figure 3 top right). The time required for inference on $1000$ testing points (Figure 3 bottom right) is an order of magnitude lower compared to the time required to train a new encoder with $500$ samples(Figure 3 bottom left).
V Simulation Data Study
@@
V-A Forward and backward transfer in a simple environment
Consider a very simple two-task environment: Gaussian XOR and Gaussian Exclusive NOR (XNOR) (Figure 4A, see Appendix A-D for details). The two tasks share the exact same discriminant boundaries: the coordinate axes. Thus, transferring from one task to the other merely requires learning a bit flip of the class labels. We sample a total 750 samples from XOR, followed by another 750 samples from XNOR.
@@ V-A Forward and backward transfer in a simple environment
V-B Forward and backward transfer for adversarial tasks
@@
VI Benchmark data Study
For benchmark data, we build SiLLy-N encoders using the network architecture described in [30]. We use the same network architecture for all benchmarking models. For the following experiments, we consider two modalities of real data: vision and language.
VI-A Reference algorithms
We compared our approaches to $16$ reference lifelong learning methods. Among them five are resource growing as well as modular approach: ProgNN [21], DF-CNN [22], LMC [25], Model Zoo[20], CoSCL [61]. Note that ”Model Zoo” was published after our work was archived on arXiv, and the authors have built on our work (personal communications). Other reference algorithms are resource constrained: Elastic Weight Consolidation (EWC) [11], Online-EWC (O-EWC) [12], Synaptic Intelligence (SI) [14], Learning without Forgetting (LwF) [15], RanDumb [62] and “None”.
We also compare two variants of exact replay (Total Replay and Partial Replay) using the code provided in [30]. Both Total and Partial Replay store all the data they have ever seen, but Total Replay replays all of it upon acquiring a new task, whereas Partial Replay replays $N$ samples, randomly sampled from the entire corpus, whenever we acquire a new task with $N$ samples. The above two replay approaches can be considered as two variants of GDumb [63]. Additionally, we have compared our approach with more constrained ways of replaying old task data, including Averaged Gradient Episodic Memory (A-GEM) [64], Experience Replay (ER) [33] and Task-based Accumulated Gradients (TAG) [65].
For the baseline “None”, the network was incrementally trained on all tasks in the standard way while always only using the data from the current task. The implementations for all of the algorithms are adapted from open source codes [22, 66]; for implementation details, see Appendix A-B.
@@ VI-A Reference algorithms
Figure 6: Pretrained encoders on CIFAR 10X10. Using pretrained encoders results in better forward transfer and accuracy for SiLLy-N.
VI-B Core benchmarks
VI-B1 CIFAR 10X10
The CIFAR 100 challenge [67], consists of 50,000 training and 10,000 test samples, each a 32x32 RGB image of a common object, from one of 100 possible classes, such as apples and bicycles. CIFAR 10x10 divides these data into 10 tasks, each with 10 classes [22] (see Appendix A-E for details).
SiLLy-N and Model Zoo demonstrate positive forward and backward transfer for every task in CIFAR 10x10, in contrast, other algorithms do not exhibit any positive backward transfer (Figure 1 first column). Moreover, they retained their accuracy while improving transfer (Figure 1 first column, first and fourth rows). ProgNN had a similar degree of forward transfer, but zero backward transfer, and requires quadratic space and time in sample size, unlike SiLLy-N which requires quasilinear space and time.
VI-B2 Spoken Digit
In this experiment, we used the Spoken Digit dataset [68]. As shown in Figure 1 second column, SiLLy-N shows positive backward and forward transfer between the spoken digit tasks, in contrast to other methods, some of which show only forward transfer, others show only backward transfer, with none showing both and some showing neither. See Appendix A-E for details of the experiment.
VI-B3 FOOD1k 50X20 Dataset
In this experiment, we use Food1k which is a large scale vision dataset consisting of $1000$ food categories from Food2k [69]. FOOD1k 50X20 splits these data into $50$ tasks with $20$ classes each. For each class, we randomly sampled $60$ samples per class for training the models and used rest of the data for testing purpose. Because on the CIFAR experiments Model Zoo performs the best among the reference resource growing models, and LwF is the best performing resource constrained algorithm, we only use them as the reference models for the large scale experiment to avoid heavy computational cost. As shown in Figure 1 third column, SiLLy-N performs the best among all the algorithms on this large dataset.
See Appendix A-E for experiments with datasets having more samples per task. In lifelong learning, we are often primarily concerned with situations in which we have a small number of samples per task. If we have enough samples per task, the learning agent does not need to transfer knowledge from other tasks. However, below we also experiment with non-trivial lifelong learning setting where sample per task is high.
These experiments indicate that SiLLy-N has positive transfer in learning environments with various classes and sample sizes.
VI-C Ablation experiments
Our proposed algorithms can improve performance on all the tasks (past and future) by both growing additional resources and replaying data from the past tasks. Below we do two ablation experiments using CIFAR 10X10 to measure the relative contribution of resource growth and replay to the performance of our proposed algorithms.
@@ VI-C Ablation experiments
These experiments indicate that (i) constraining the resource growth results in lower accuracy while still achieves positive forward and backward transfer for the algorithm, (ii) lowering the amount of replay lowers backward transfer without any effect on forward transfer.
Experiment using pretrained encoders
We explore the effect of using pretrained encoders on the performance of SiLLy-N. For this experiment only, we use ResNet 50 and vision transformer ViT_B16 (provided in keras-vit package) [70]. We freeze all the layers excluding the final two linear layers during training. Pretraining the encoders results in better accuracy and forward transfer, but less backward transfer for SiLLy-N (Figure 6). Vision transformers achieved lower accuracy, as expected with the small samples sizes used here [71]. Pretrained network requires less training epochs for each task (with early stopping). In this experiment, one forward pass through ViT takes $\sim 5$ seconds and training a ViT encoder ($20$ epochs) takes $\sim 4$ minutes using an Apple M1 Max chip and $64$ GB of RAM.
VI-D Adversarial analysis
@@
Figure 7: Adversarial CIFAR 10x10 experiments. Left: Permuting the task sequenc does not affect $\mathsf{Transfer}$ for SiLLy-N. The mean position, indicated by the black dash, and the spread of the dots on both sides of the mean remain nearly the same. Middle: Shuffling class labels within tasks two through nine with 500 samples each demonstrates SiLLy-N can still achieve positive backward transfer, and that the other algorithms still fail to transfer. Right: SiLLy-N is nearly invariant to rotations, whereas other approaches are more sensitive to rotation.
Consider the same CIFAR 10x10 experiment setup in Section VI-B1. In the following experiments, we modify the above setup in different adversarial settings.
Shuffled task order experiment
We repeat the same experiment mentioned above $5$ times more by permuting the task order. Figure 7 left column shows the mean as well as the spread of $\mathsf{Transfer}$ over all tasks remain the same for different shuffled task orders. Note that the encoders and thus all the channels remain the same regardless of the above task order permutation after all the tasks have been introduced, hence the distribution of $\mathsf{Transfer}$ for all the tasks after all the tasks have been introduced remain similar.
Label shuffle experiment
In this experiment, for Task 2 through 10, randomly permute the class labels within each task, rendering each of those tasks adversarial with respect to the first task (because the labels are uninformative). Figure 7 middle column indicates that SiLLy-N show positive backward transfer even with such label shuffling (the other algorithms, except Model Zoo, did not demonstrate positive backward transfer).
Rotation experiment
Consider a Rotated CIFAR experiment, which uses only data from the first task, divided into two subsets of equal size (making two tasks), where the second subset is rotated by different amounts (Figure 7 right column). Backward transfer of SiLLy-N is nearly invariant to rotation angle, whereas the other approaches are much more sensitive to the rotation angle. Note that the zero rotation angle corresponds to the two tasks having identical distributions. The fact that other algorithms fail to transfer even in this setting suggests that they may never be able to achieve a positive backward transfer. See Appendix A-E for an additional experiment using CIFAR 10X10.
These adversarial experiments indicate that SiLLy-N is robust to adversarial perturbations in the source tasks, while most of the other algorithms are not.
VI-E Constant Resource Mode Operation